# An active learning method for solving competitive multi-agent decision-making and control problems

### v4

@@ IV-D Asymptotic properties
As a consequence of Theorem 4.5, in case the convergence of the parametric estimates happens, it can only be towards the true (at least pointwise) values. This fact, together with $\lim_{k\to\infty}\|\hat{\boldsymbol{x}}^{k}-\boldsymbol{x}^{k}\|_{2}=0$, yields $\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-f_{i}(\hat{\boldsymbol{x}}_{-i})\|^{2}_{2}=0$, and hence $\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|_{2}=0$ for all $i\in\mathcal{N}$, i.e., the external entity has in its hands both a stationary action profile $\hat{\boldsymbol{x}}$, and pointwise-exact surrogates of the action-reaction mappings $\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})=f_{i}(\hat{\boldsymbol{x}}_{-i})$.
@@ A-A Technical proofs
In view of the definition of limit we can then always find some $\bar{k}\in\mathbb{N}$ and $\mu_{i}>0$ such that a neighbourhood of $\bar{x}_{i}$, say $\mathcal{B}_{\mu_{i}}(\bar{x}_{i})$, contains infinitely many points [53], meaning that $x_{i}^{k}\in\mathcal{B}_{\mu_{i}}(\bar{x}_{i})$ for all $k\geq\bar{k}$, $i\in\mathcal{N}$. By virtue of the consistency property proved in Lemma 4.2 it thus follows that the pointwise approximation shall be exact, namely each $\tilde{\theta}_{i}$ is so that $\|\hat{f}_{i}(\bar{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})-f_{i}(\bar{\boldsymbol{x}}_{-i})\|_{2}=0$, i.e., $\tilde{\theta}_{i}\in\mathcal{A}_{i}=\mathcal{A}_{i}(\tilde{x}_{i},\bar{\boldsymbol{x}}_{i})$.
This latter relation has hence two main consequences. First, it excludes the possibility that, due to the presence of some tie-break rule $\mathcal{T}(\cdot)$, $\lim_{k\to\infty}\|\hat{\boldsymbol{x}}^{k}-\boldsymbol{x}^{k}\|_{2}\neq 0$, namely the external entity can not learn any map but the true one locally, thus forcing $\tilde{\boldsymbol{x}}=\bar{\boldsymbol{x}}$. If this were not the case, indeed, the update rule for each $\theta_{i}$ would be fed with different samples, thus possibly preventing their convergence (this would also happen if a tie-break rule for singling out a minimizer from (6) were not adopted). Furthermore, in view of Definition 2.1 and since (7) is solved at the global optimum at every $k\in\mathbb{N}$, it shall happen that $\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-\hat{f}_{i}(\hat{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\hat{\boldsymbol{x}}_{i}-f_{i}(\hat{\boldsymbol{x}}_{-i})\|^{2}_{2}=0=\sum_{i\in\mathcal{N}}\|\bar{\boldsymbol{x}}_{i}-f_{i}(\bar{\boldsymbol{x}}_{-i})\|^{2}_{2}$, and hence $\|\bar{\boldsymbol{x}}_{i}-f_{i}(\bar{\boldsymbol{x}}_{-i})\|_{2}=0$ for all $i\in\mathcal{N}$. This latter $\bar{\boldsymbol{x}}\in\Omega$ therefore coincides with a fixed point of the action-reaction mappings $f_{i}(\cdot)$, and hence with a stationary action profile in the sense of Definition 2.1, concluding the proof. $\blacksquare$

### v5

@@ IV-D Asymptotic properties
As a consequence of Theorem 4.5, in case the convergence of the parametric estimates happens, it can only be towards the true (at least pointwise) values. This fact, together with $\lim_{k\to\infty}\|\hat{\boldsymbol{x}}^{k}-\boldsymbol{x}^{k}\|_{2}=0$, yields $\sum_{i\in\mathcal{N}}\|\tilde{x}_{i}-\hat{f}_{i}(\tilde{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\tilde{x}_{i}-f_{i}(\tilde{\boldsymbol{x}}_{-i})\|^{2}_{2}=0=\sum_{i\in\mathcal{N}}\|f_{i}(\bar{\boldsymbol{x}}_{-i})-\bar{x}_{i}\|^{2}_{2}=0$, i.e., $\|f_{i}(\bar{\boldsymbol{x}}_{-i})-\bar{x}_{i}\|_{2}=0$, i.e., the external entity has in its hands both a stationary action profile $\tilde{\boldsymbol{x}}$, and pointwise-exact surrogates of the action-reaction mappings $\hat{f}_{i}(\tilde{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})=f_{i}(\tilde{\boldsymbol{x}}_{-i})$.
@@ A-A Technical proofs
In view of the definition of limit we can then always find some $\bar{k}\in\mathbb{N}$ and $\mu_{i}>0$ such that a neighbourhood of $\bar{x}_{i}$, say $\mathcal{B}_{\mu_{i}}(\bar{x}_{i})$, contains infinitely many points [53], meaning that $x_{i}^{k}\in\mathcal{B}_{\mu_{i}}(\bar{x}_{i})$ for all $k\geq\bar{k}$, $i\in\mathcal{N}$. By virtue of the consistency property proved in Lemma 4.2 it thus follows that the pointwise approximation shall be exact, namely each $\tilde{\theta}_{i}$ is so that, for all $i\in\mathcal{N}$, $\|\hat{f}_{i}(\tilde{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})-f_{i}(\tilde{\boldsymbol{x}}_{-i})\|_{2}=0=\|\hat{f}_{i}(\tilde{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})-\bar{x}_{i}\|_{2}$.
Since (7) is solved at the global optimum at every $k\in\mathbb{N}$, and the single-valuedness of each $f_{i}(\cdot)$, this latter relation readily yields $\sum_{i\in\mathcal{N}}\|\tilde{x}_{i}-\hat{f}_{i}(\tilde{\boldsymbol{x}}_{-i},\tilde{\theta}_{i})\|^{2}_{2}=\sum_{i\in\mathcal{N}}\|\tilde{x}_{i}-f_{i}(\tilde{\boldsymbol{x}}_{-i})\|^{2}_{2}=0=\sum_{i\in\mathcal{N}}\|\tilde{x}_{i}-\bar{x}_{i}\|^{2}_{2}$, and hence $\|\tilde{x}_{i}-\bar{x}_{i}\|_{2}=0$ for all $i\in\mathcal{N}$. This means that $\bar{\boldsymbol{x}}=\tilde{\boldsymbol{x}}$, and in view of Definition 2.1, they coincide with a fixed point of the action-reaction mappings $f_{i}(\cdot)$, as $\sum_{i\in\mathcal{N}}\|f_{i}(\bar{\boldsymbol{x}}_{-i})-\bar{x}_{i}\|^{2}_{2}=0$, i.e., $\|f_{i}(\bar{\boldsymbol{x}}_{-i})-\bar{x}_{i}\|_{2}=0$ for all $i\in\mathcal{N}$, concluding the proof. $\blacksquare$
@@ A-B Derivation of Kalman filter equations in (12)
Acknowledgment
The authors thank Prof. Barbara Franci, Maastricht University, for the fruitful discussions on the technical proofs.